OpenStudy (anonymous):

Can someone help me solve these word problems? The sum of two numbers is 28 and their differences is 6. Find each of the numbers.

OpenStudy (noorik.):

you would suppose the first number x and the second number y so x+y=28 x-y=6

OpenStudy (anonymous):

Idk how to do these

OpenStudy (noorik.):

try to sum the two equations like (x+y) + (x-y) = 28 + 6 2x = 34 x = ? then substitute x in the first equation to find y

OpenStudy (anonymous):

I got it thanks

OpenStudy (noorik.):

you are welcome